A hospital director is told that 54% of the treated patients are insured. The director wants to test the claim that the percentage of insured patients is less than the expected percentage. A sample of 50 patients found that 75 were insured. Find the value of the test statistic. Round your answer to two decimal places.
Answer #1

Solution :

= x / n = 50 / 75 = 0.6667

P0 = 0.54

1 - P0 = 0.46

Test statistic = z

= - P0 / [P0 * (1 - P0 ) / n]

= 0.6667 - 0.54 / [(0.54 * 0.46) / 75]

= 2.20

Test statistic = z = 2.20